CURRICULUM VITA

Huaiqing Zuo

CONTACT INFO

Department of Mathematics, Statistics, & Computer Science 322 Science & Engineering Offices (M/C 249) 851 S. Morgan Street Chicago, IL 60607-7045

phones: (773)-289-6831. e-mail: hqzuo@126.com, hqzuo@math.uic.edu

INTERESTS

Algebraic Geometry and Toric Code

EDUCATION

2007-2009 Exchange visiting student in Mathematics, University of Illinois at Chicago, and PH.D. Graduate Student of ECNU.

USA; Thesis: ”TBA”, advisor Stephen. Yau (May 2009).

2004-2007 M.S. in Mathematics , East China Normal University, Shanghai, P. R. China, advisor Zhijie Chen & Shengli tan.

1998-2004 B.S. in Mathematics, Guangzhou University.

APPOINTMENTS

2002,7-2004,9 Lecturer at Guangzhou No.6 high school.

List of Publications

1. Notes on classification of toric surface codes of dimension 5 (with Stephen S.-T. Yau), Applicable Algebra in Engineering, Communication and Computing, 20 (2009), 175-185.

2. Classification of Gradient Space of Dimension 8 With a Reducible sl(2,ℂ) action (with Yung Yu and Stephen S.-T. Yau) 44pp. Science in China Series A: Mathematics 2009, Vol. 52, No. 1-35.

3. ON THE PARAMETERS OF TWO SPECIAL TORIC SURFACE CODES (with Ping Zhang), 9 pp. To appear at .Chinese Annals of Mathematics,Series A and C.

4. Lojasiewicz Inequality For Weighted Homogeneous Polynomial With Isolated Singularities (with Shengli Tan and Stephen S.-T. Yau), 12pp. To appear at Proceeding of AMS .

5. Biggest Sharp Polynomial Estimate of Integral Points in Rightangled Simplices (with Linda Zhao and Stephen S.-T. Yau), 43 pp. To appear at Contemporary Mathematics.

TALKS

1. K3 surface and Torelli theorem, Algebra Geometry seminar, ECNU, spring, 2007

2. Normal 2 dimensional singularities, Graduate Student Algebra Geometry seminar, University of Illinois at Chicago, spring, 2008

3. Introduction to toric variety, The UIC Algebraic Geometry Seminar , University of Illinois at Chicago, spring, 2008

4. Moduli of Curve, Graduate Student Algebra Geometry seminar, University of Illinois at Chicago, fall, 2008

CONFERENCES

1. Conference on algebraic Geometry, Fu dan university September 14 - September 16, 2005

2. Forum about Chinese Mathematic and Eduction, Zhejiang university, Hangzhou 6.30-7.2 2006

3. Summer school at East China Normal University, summer, 2006 4. The Second Chinese-German Conference ” Komplexe Geometrie

(Complex Geometry)” East China Normal University, Shanghai, China September 11 - September 16, 2006

5. The 10th National Conference on Algebra, Xiamen October 22 - October 27, 2006

6. Summer school at East China Normal University, summer, 2007 7. Midwest Number Theory Days, University of Illinois at Chicago,

March 7-8 2008

8. A Conference to celebrate Robin Hartshorne’s 70th birthday, University of Illinois at Chicago, April 11-13 2008

9. 2nd Bluegrass Algebra Conference, Department of Mathematics University of Kentucky Lexington, March 6-8, 2009

AWARDS

1. Excellent graduate student scholarship at East China Normal University 2005

TEACHING EXPERIENCE

1. College Algebra.

2. Calculus.